# A DNN Biophysics Model with Topological and Electrostatic Features

### v2

@@ 2.3.1 Simplicial Homology
$\sigma^{k}=\left\{\sum_{i=0}^{k}\lambda_{i}v_{i}\;\middle|\;\lambda_{i}\geq 0,\sum\lambda_{i}=1\right\}$
@@
$\partial_{k}[v_{0},v_{1},\dots,v_{k}]=\sum_{i=0}^{k}(-1)^{i}[v_{0},\dots,\hat{v}_{i},\dots,v_{k}]$
@@
With the concepts of simplex, face, boundary, chain, etc., we can formally define that a simplicial complex $K$ is a collection of simplices such that:
@@
Thus, each simplicial complex corresponds to a topological space by “filling in” the simplexes and gluing them together along shared faces.
2.3.2 Persistent Homology
However, the simplicial homology of a complex associated to a point set at a particular proximity parameter $\epsilon$ is insufficient to characterize the signal or noise. This calls for the ideas of filtration and persistent homology, which can identify and connect complexes at different level of complexity and record the appearance and disappearance of the homology group. The importance of persistence is that given a parameterized family of spaces, topological features which persist over a significant parameter ranges are considered as signal while those with short-lived features are treated as noise [53]. To this end, we introduce the concept of filtration, which is a sequence of nested simplicial complexes:
$\emptyset=K_{0}\subseteq K_{1}\subseteq\cdots\subseteq K_{n}=K.$
Each $K_{i}$ is a simplicial complex, and the sequence grows over time or scale. Instead of computing just the $k$-homology group $H_{k}(K)$, which is an algebraic object capturing the $k$-dimensional holes in a topological space represented by a simplicial complex $K$, we compute $H_{k}(K_{i})$ for each $i$ and track features as the birth, death, and persistence of a homology class. Let $C_{k}(K)$ be the group of $k$-chains, and let
$\cdots\xrightarrow{\partial_{k+1}}C_{k}(K)\xrightarrow{\partial_{k}}C_{k-1}(K)\xrightarrow{\partial_{k-1}}\cdots$
be the sequence of boundary maps, the $k$-homology group of $K$ is defined as:
$H_{k}(K)=\frac{\ker(\partial_{k})}{\operatorname{im}(\partial_{k+1})}$
@@ 2.3.3 Feature Generation from Barcode
To make the topological features uniform and physically informed, we choose to use Element Specific Persistent Homology (ESPH) to extract topological features at different level of complexity. ESPH has been used for successfully predicting protein-ligand binding [19]. The main idea is to select atom elements, which form the cloud of points, using target problem related physical, chemical, and biological information. In this project, since solvation energy and Coulomb energy are highly related to the amino acid chains, we use two collections of cloud of points. One is the collection of all Carbon atoms, which form the backbone of amino acids. The other is the collection of all heavy atoms {C, N, O, S}, whose locations and interconnections are also critical in protein structures. Once the collections of cloud of points are determined and selected, the barcode [53] carrying persisting homology information of the cloud of points are produced. Here the barcode is a graphical representation of persistent homology as a collection of horizontal line segments in a plane whose horizontal axis corresponds to the proximity parameter and whose vertical axis represents an (arbitrary) ordering of homology generators such as $H_{0}$, $H_{1}$, $H_{2}$, etc. We use the software GUDHI [54] to implement this task. The produced barcode is then used by the algorithm below to generate the vector of topological features.
Figure 2: The barcode generated by the collection of cloud points using heavy atoms in proteins (a) 1a1e and (b) 1pb9.
Figure 2 gives the barcode for protein 1a1e, the C-SRC Tyrosine Kinase, and protein 1pb9, the N-methyl-D aspartate receptor, using heavy atoms. The persistent homology analysis conducted in this study characterizes key biophysical phenomena, including intermolecular interactions, solvation effects, and hydrophobicity. Topological features are extracted from Betti-1 and Betti-2 barcodes, where the birth, death, and persistence values are computed based on the spatial distribution of carbon atoms and all heavy atoms within the protein structure, resulting in a total of 12 feature channels. Features derived from all heavy atoms represent the geometric interactions between specific types of atoms in the protein, whereas those computed from carbon atoms primarily reflect hydrophobic interactions and indirectly capture solvation effects. The analysis is performed over a distance scale of 0 to 50Å, which is discretized into uniform bins of 0.50Å to ensure fine-grained resolution of topological patterns. The topological information included in the barcode can be converted to features with the following approach [19].
@@
$\mathcal{D}$ : dimension (i.e., Betti-0, Betti-1, etc.)
Using the collection, the structured vectors ${\bf V}^{b}$, ${\bf V}^{d}$, and ${\bf V}^{p}$ can be constructed to respectively describe the birth, death, and persistent patterns of the barcodes in various spatial dimensions. Practically, the filtration interval $[0,L]$ is divided into $n$ equal length subintervals and the patterns are characterized on each subinterval. The description vectors using their $i$th component are defined as:
@@ 2.4.1 Introduction
In most of the molecular simulations, including Monte Carlo simulation, Brownian Dynamics, Molecular Dynamics, etc., the electrostatic interactions are characterized by the partial charges assigned at the atomic centers. These partial charges are assigned according to the force fields, which are a set of mathematical equations and parameters used to calculate the potential energy of a molecule based on its atomic structure and interactions. The force fields are determined by experiment or quantum chemistry.
Figure 3: An 2-d illustration for 3-d uniform and multi-scale electrostatic features for a protein (purple dashed line) with charges (shown as black dots); left: charges $q$ and reaction potential $\phi_{\text{reac}}$ are redistributed as point-multipoles (shown as explosion symbols) using Cartesian treecode [45] or FMM [46] at the centers of the cluster at different levels (level 0: black; level 1: red; level 2: green); right: charges $q$ and reaction potential $\phi_{\text{reac}}$ are redistributed at Chebyshev nodes (shown as square: level 0: black; level 1: red; level 2: green; only one cluster is shown in each level) using Barycentric Lagrangian treecode [56].
The algorithm for obtaining the mulit-scale, physics-informed, uniform electrostatic features can be explained using the left of Fig. 3. The electrostatic profile of the protein is determined by the partial charges $q_{n}({\bf r}_{n})$ or reaction potential $\phi_{\text{reac}}({\bf r}_{n})$ at the atomic centers ${\bf r}_{n}$’s as illustrated by the black dots for $n=1,2,\cdots,N_{c}$ with $N_{c}$ as the number of atoms of the protein. Partial charges are the source of electrostatic interactions while reaction potentials are the outcomes of electrostatic interactions for the solvated potein, which includes the important solute-solvent interactions. Consider the variation of $N_{c}$, the number of atoms, from protein to protein, it is difficult to directly use $q_{n}({\bf r}_{n})$ and/or $\phi_{\text{reac}}({\bf r}_{n})$ as features for machine learning. Our strategy to alleviate this pain is to use point-multipoles as illustrated by the explosion symbols at the cluster centers to carry redistributed charges or reaction potentials in terms of moments of the multipole expansion. The number of cluster centers are uniform for proteins with various numbers of atoms. It is also multiscale as being determined by the combination of levels of the tree $L$ and number of terms of the multipole expansion $p$. The users specify $L$ and $p$ as needed.
In the left of Fig. 3. the $N_{d}$ point multipoles (explosion symbols) are organized in a hierarchical way such that different level of details are given by different number of symbols at different levels. For the $m$th multipole, $m=1,\cdots,N_{d}$, using a 3-d index ${\bf k}=(k_{1},k_{2},k_{3})$, $|{\bf k}|=k_{1}+k_{2}+k_{3}$, in a cluster $c$ with cluster center ${\bf x}_{c}$, the moments are defined as
@@
If the Cartesian treecode algorithm [45] is used, the computational cost is $O(N_{c}\log N_{c})$ for obtaining all the moments as detailed in [45], where $N_{c}$ is number of the particles/charges and $\log N_{c}$ is the level of trees. In fact, the computational cost can be as low as $O(N_{c})$ using the strategies in [46] such that the moments at the finest cluster are computed first and a M2M (moments to moments) transformation can be used to efficiently compute moments at any desired level from the bottom to the top. Our simulations so far showed that the $O(N_{c}\log N_{c})$ cost is sufficiently fast for moments computation since $N_{c}$ is within the scale of tens of thousands even for large proteins. These moments are intrinsic properties of the cluster thus can serve as features for the protein, which carries simplified but important structural and charge information.
@@
Alternatively, the Barycentric treecode approach as shown in the right picture of Fig. 3 can be used to generate the electrostatic features located the Chebyshev points (the squares) with modified weights as feature values. The details can be found in [56]. For this approach, the computational cost is $O(N_{c}\log N_{c})$, and the number of terms in the similar sense of $N_{p}$ is $O(c_{p}^{3})$ where $c_{p}$ is the number of Chebyshev nodes in one direction. In this project, we focus on the Cartesian treecode approach, while the Barycentric treecode approach will be investigated in the future.
@@ 2.4.2 Implementation details
Step 1. The protein structure (e.g. the PDB file) is obtained from publicly available repo e.g. the protein data bank [5]. The protein is represented by its atomic locations and partial charges: $q_{i}({\bf r}_{i})$ for $i=1,\cdots,N_{c}$, e.g. the PQR file using the PDB2PQR software [57].
@@
Step 3. The Cartesian treecode [45] will be called and the moments for the multipoles will be output as the electrostatic features. These features can be used as part of the inputs for machine learning algorithms, which will be shown the next.
2.5 A DNN based machine learning model with topological and electrostatic features
@@
Figure 4: The DNN based Machine Learning model which uses protein structural data, force field, and known protein properties to mathematically generate electrostatic, topological, algebraic, geometric, and features to train a learned model and then use the learned model to predict unknown properties for protein with available structures.
@@
The main goal of this project is to discover the hidden and useful information embedded in the protein structural data from the protein data bank in a simple and abstract way. The protein structure data and force field include bonded and non-bonded interactions with which the molecular simulation can be performed. We categorize these information into electrostatic, topological, algebraic, and geometric features. These four aspects are all important to the machine learning model. In this project, we put our focus on the electrostatic and topological features. The DNN based Machine Learning model is shown in Fig. 4, which uses the available protein structural data [58], force field such as AMBER [59], CHARMM[60], AMOEBA [61], etc. and known protein properties repositories such as PDBbind Database [62], Protein pKa Database [63], etc to mathematically generate features to train a learned model and then use the learned model to predict unknown properties for protein with available structures. In this project, we focus on the effectiveness of topological and electrostatic features using computed Coulombic energies and solvation free energies as labels.

### v3

@@ 2.3.1 Simplicial Homology
$\sigma^{k}=\left\{\sum_{i=0}^{k}\lambda_{i}v_{i}\;\middle|\;\lambda_{i}\geq 0,\sum\lambda_{i}=1\right\}$ (4)
@@
$\partial_{k}[v_{0},v_{1},\dots,v_{k}]=\sum_{i=0}^{k}(-1)^{i}[v_{0},\dots,\hat{v}_{i},\dots,v_{k}]$ (5)
@@
With the concepts of simplex, face, boundary, chain, etc., we define that a simplicial complex $K$ is a collection of simplices such that:
@@
Thus, each simplicial complex corresponds to a topological space by filling in the simplexes and gluing them together along shared faces.
Let $C_{k}(K)$ be the group of $k$-chains, and let
$\cdots\xrightarrow{\partial_{k+1}}C_{k}(K)\xrightarrow{\partial_{k}}C_{k-1}(K)\xrightarrow{\partial_{k-1}}\cdots$ (6)
be the sequence of boundary maps $\partial_{k}$, the $k$-homology group of $K$ is defined as:
$H_{k}(K)=\frac{\ker(\partial_{k})}{\operatorname{im}(\partial_{k+1})}$ (7)
@@
2.3.2 Persistent Homology
However, the simplicial homology of a complex associated to a point set at a particular proximity parameter $\epsilon$ is insufficient to characterize the signal or noise. This calls for the ideas of filtration and persistent homology edelsbrunner:2000, zomorodian:2005, which can identify and connect complexes at different level of complexity and record the appearance and disappearance of the homology group. To this end, we introduce the concept of filtration, which is a sequence of nested simplicial complexes:
$\emptyset=K_{0}\subseteq K_{1}\subseteq\cdots\subseteq K_{n}=K.$ (9)
Each $K_{i}$ is a simplicial complex, and the sequence grows over time or scale. Instead of computing just the $k$-homology group $H_{k}(K)$, which is an algebraic object capturing the $k$-dimensional holes in a topological space represented by a simplicial complex $K$, we compute $H_{k}(K_{i})$ for each $i$, the index for time or scale parameters.
As $i$ or its corresponding scale parameter grows, the duration, as measured by the change of $i$ or its corresponding scale parameter, of the homology class $H_{k}(K_{i})$ between its appearance and disappearance is called persistence. The importance of persistence is that given a parameterized family of spaces, topological features which persist over a significant parameter ranges are considered as signal while those with short-lived features are treated as noise Ghrist:2007. In practice, the persistent homology is computed using the pipeline that involves steps of turning data into a sequence of spaces (a filtration), tracking the birth and death of topological features using algebra, implementing using linear algebra tools, and visualization. The pipeline is realized by many established software, e.g. GUDHI Boissonnat:2014, as our choice.
We next introduce the barcode, persistent diagram, and persistent image as approaches to visualize the persistent homology. These approaches are different in format and visualization but are equivalent mathematically. Following the definition, we provide the algorithms to extract topological features from barcode and persistent image.
2.3.3 Barcode, Persistence Diagram and Persistence Image
Persistent homology captures the evolution of topological features throughout a filtration process, which could be encoded in the forms of barcode, persistent diagram, and persistent image.
A barcode for dimension $k$ is the multiset $\{[b_{j},d_{j})\}_{j}$, where each interval corresponds to a $k$-dimensional homology class with index $j$, which appears at $b_{j}$ and disappears at $d_{j}$. In the picture of a barcode, each class is visualized as a bar starting at $b_{j}$ and ending at $d_{j}$. Thus the barcode is a graphical representation of persistent homology as a collection of horizontal line segments in a plane whose horizontal axis corresponds to the filtrition parameter and whose vertical axis represents an (arbitrary) ordering of homology generators such as $H_{0}$, $H_{1}$, $H_{2}$, etc. Different colors can be used to distinguish different dimensions.
A persistent diagram represents each homology class as a point $(b_{j},d_{j})$ in the birth–death plane, where $b_{j}$ and $d_{j}$ denote the filtration parameters at which the class appears and disappears. The persistent diagram thus is a scatter plot of points, where distance from the diagonal measures their significance.
From their definitions, we can tell barcode and persistence diagram carry equal amount of information. We therefore only use barcode to generate topological features. The persistent image, which is built on top of persistent diagram to generate fixed-size 2-d image, is readily used as features or inputs for the convolution neural network. Details about persistent image and its formal mathematical definition can be found in supplementary materials.
@@ 2.3.4 Feature generation from Barcode
To make the topological features uniform and physically informed, we choose to use Element-Specific Persistent Homology (ESPH) to extract topological features at different levels of complexity. ESPH has been successfully applied to the prediction of protein–ligand binding. Cang:2017. The core idea is to select atomic elements that form the point cloud by incorporating physical, chemical, and biological information relevant to the target problem. In this project, since solvation energy and Coulomb energy are closely related to the amino acid chains, we use two collections of point clouds. One consists of all carbon atoms, which form the backbone of amino acids. The other consists of all heavy atoms {C, N, O, S}, whose locations and interconnections are also critical to protein structures. Once the collections of point clouds are determined, the corresponding barcodes Ghrist:2007, which encode the persistent homology of the point clouds, are constructed. We use the GUDHI software package Boissonnat:2014 to generate these barcodes. The resulting barcodes are then processed by the algorithm below to produce vectors of topological features.
Figure 2: The barcode of $H_{0},H_{1},H_{2}$ generated by the collection of cloud points using heavy atoms in proteins (a) 1a1e and (b) 1pb9.
Figure 2 shows the barcodes of $H_{0},H_{1},H_{2}$ for protein 1a1e, the C-SRC Tyrosine Kinase, and protein 1pb9, the N-methyl-D aspartate receptor, using heavy atoms. The persistent homology analysis conducted in this study characterizes key biophysical phenomena, including intermolecular interactions, solvation effects, and hydrophobicity. Topological features are extracted from $H_{1}$ and $H_{2}$ barcodes, where the birth, death, and persistence values are computed based on the spatial distribution of carbon atoms and all heavy atoms within the protein structure, resulting in a total of 12 feature channels.
The significance of 1- and 2-dimensional homology groups lies in the fact that they represent 2-d rings and 3-d voids, respectively, which persist over a wide range of scales and are both topologically and geometrically meaningful. In contrast, the 0-dimensional homology group represents connected components, whose number decreases from the total number of points to 1 and therefore carries limited persistent signal. Higher-dimensional homology groups, although mathematically defined, are typically short-lived in persistence, behave as combinatorial noise, and lack clear geometric meaning. We use both the collection of heavy atoms and the collection of carbon atoms to capture critical structural and chemical information of proteins. Features derived from all heavy atoms represent geometric interactions between specific types of atoms in the protein, whereas those computed from carbon atoms primarily reflect hydrophobic interactions and indirectly capture solvation effects. The analysis is performed over a distance scale of 0 to 50Å, which is discretized into uniform bins of 0.25Å  to ensure fine-grained resolution of topological patterns. The topological information encoded in the barcodes is then converted into features using the following approach Cang:2017.
@@
$\mathcal{D}$ : dimension (i.e., $H_{1}$, $H_{2}$, etc.)
Using the collection, the structured vectors ${\bf V}^{b}$, ${\bf V}^{d}$, and ${\bf V}^{p}$ can be constructed to respectively describe the birth, death, and persistent patterns of the barcodes in various spatial dimensions. Practically, the filtration interval $[0,L]$ is divided into $n$ equal length subintervals and the patterns are characterized on each subinterval. The description vectors using their $i$th component for $1\leq i<n$, are defined as:
@@
2.3.5 Feature generation from Persistent Image
The topological features can also be constructed using persistence image alternatively. To this end, we first extract persistence diagrams from protein structures and convert them into stable, fixed-size image representations. The persistence diagrams are obtained separately from two types of point clouds: the collection of all carbon atoms and the collection of all heavy atoms. These two approaches are also used in the barcode-based feature construction as detailed previously. For each approach, we consider topological features in homology groups $H_{1}$ and $H_{2}$. Each persistence diagram is filtered using physically meaningful thresholds such that birth and persistence values are restricted to maximum bounds of $12.5$ Å. The resulting persistence domain is discretized using a resolution of $0.25$ Å, which produces a uniform image grid $D=[0,b_{\text{thr}}]\times[0,p_{\text{thr}}]$ with $b_{\text{thr}}=p_{\text{thr}}=50$ for all proteins. To emphasize the contribution of long-lived topological features, each point $(b_{i},p_{i})\in D$ is assigned a smooth exponential weight
$\omega(b_{i},p_{i})=e^{p_{i}}-1.$ (11)
(a) 1a0t (b) 1a4q
Figure 3: Barcode (top) , persistence diagrams (middle) and persistence image (bottom) for proteins 1a0t (a) and 1a4q (b). For each protein, from left to right are: $H_{1}$ and $H_{2}$ for carbon atoms, $H_{1}$ and $H_{2}$ for all heavy atoms.
This exponential weighting enhances features with larger persistence while assigning near-zero weight to short-lived topological noise, thereby improving the contrast between significant and insignificant structures in the resulting persistence images. In this project, persistence images are generated using the Gaussian kernel with bandwidth of $0.4$ Å, which is implemented by using the GUDHI library Boissonnat:2014. For each protein, four persistence images are produced: carbon atoms only for $H_{1}$ and $H_{2}$, all heavy atoms for $H_{1}$ and $H_{2}$. These four images are stacked to form a four-channel topological feature tensor, $\mathbf{X}\in\mathbb{R}^{b_{\text{thr}}\times p_{\text{thr}}\times 4}$. Finally, the resulting persistence image tensor is stored in the NumPy format and used as the input topological representation for downstream deep learning models. This approach provides a stable, compact, and fully vectorized encoding of protein topology thus is well suited for ML-based prediction.
In Fig. 3, the barcode of $H_{1}$ and $H_{2}$ using carbon atoms and using all heavy atoms for protein 1a0t and 1a4q are provided for comparison purpose with the persistent diagram and persist image. The persistence diagrams of the two proteins 1a0t and 1a4q show that most of the topological features are short-lived and concentrated near the diagonal, while only a small number of them persist. The corresponding persistence images highlight these persistent structures as bright and localized regions, with background noise effectively suppressed. The observed differences between persistence images from the cloud of points of carbon atoms and those from all heavy atoms show that topological invariants are influenced by the underlying chemical composition.
@@ 2.4.1 Introduction
Figure 4: An 2-d illustration for 3-d uniform and multi-scale electrostatic features for a protein (purple dashed line) with charges (black dots); Charges $q$ and reaction potential $\phi_{\text{reac}}$ are redistributed as point-multipoles (explosion symbols) using Cartesian treecode Li:2009 or FMM Tausch:2003 at the centers of the cluster at different levels (level 0: black; level 1: red; level 2: green).
In most molecular simulations, including Monte Carlo simulation, Brownian Dynamics, Molecular Dynamics, etc., the electrostatic interactions are characterized by the partial charges assigned at the atomic centers. These partial charges are assigned according to the force fields, which are a set of mathematical equations and parameters used to calculate the potential energy of a molecule based on its atomic structure and interactions. The force fields are determined by experiment or quantum chemistry.
The algorithm for obtaining the mulit-scale, physics-informed, uniform electrostatic features can be explained using Fig. 4. The electrostatic profile of the protein is determined by the partial charges $q_{n}({\bf r}_{n})$ or reaction potential $\phi_{\text{reac}}({\bf r}_{n})$ at the atomic centers ${\bf r}_{n}$’s as illustrated by the black dots for $n=1,2,\cdots,N_{c}$ with $N_{c}$ as the number of atoms of the protein. Partial charges are the source of electrostatic interactions while reaction potentials are the outcomes of electrostatic interactions for the solvated potein, which includes the important solute-solvent interactions. Consider the variation of $N_{c}$, the number of atoms, from protein to protein, it is difficult to directly use $q_{n}({\bf r}_{n})$ and/or $\phi_{\text{reac}}({\bf r}_{n})$ as features for machine learning. Our strategy to alleviate this pain is to use point-multipoles as illustrated by the explosion symbols at the cluster centers to carry redistributed charges or reaction potentials in terms of moments of the multipole expansion. The number of cluster centers are uniform for proteins with various numbers of atoms. It is also multiscale as being determined by the combination of levels of the tree $L$ and number of terms of the multipole expansion $p$. The users specify $L$ and $p$ as needed.
In Fig. 4. the $N_{d}$ point multipoles (explosion symbols) are organized in a hierarchical way such that different level of details are given by different number of symbols at different levels. For the $m$th multipole, $m=1,\cdots,N_{d}$, using a 3-d index ${\bf k}=(k_{1},k_{2},k_{3})$, $|{\bf k}|=k_{1}+k_{2}+k_{3}$, in a cluster $c$ with cluster center ${\bf x}_{c}$, the moments are defined as
@@
If the Cartesian treecode algorithm Li:2009 is used, the computational cost is $O(N_{c}\log N_{c})$ for obtaining all the moments as detailed in Li:2009, where $N_{c}$ is number of the particles/charges and $\log N_{c}$ is the level of trees. In fact, the computational cost can be as low as $O(N_{c})$ using the strategies in Tausch:2003 such that the moments at the finest cluster are computed first and a M2M (moments to moments) transformation can be used to efficiently compute moments at any desired level from the bottom to the top. Our simulations so far showed that the $O(N_{c}\log N_{c})$ cost is sufficiently fast for moments computation since $N_{c}$ is within the scale of tens of thousands even for large proteins. These moments are intrinsic properties of the cluster thus can serve as features for the protein, which carries simplified but important structural and charge information.
@@ 2.4.2 Implementation Details
Step 1. The protein structure (e.g. the PDB file) is obtained from publicly available repo e.g. the protein data bank Berman:2000. The protein is represented by its atomic locations and partial charges: $q_{i}({\bf r}_{i})$ for $i=1,\cdots,N_{c}$, e.g. the PQR file using the PDB2PQR software Dolinsky:2007.
@@
Step 3. The Cartesian treecode Li:2009 will be called and the moments for the multipoles will be output as the electrostatic features. These features can be used as part of the inputs for ML algorithms, which will be shown next.
2.5 The Machine Learning Model
@@
Figure 5: The DNN based Machine Learning model using protein structures, force field, and known protein properties to mathematically generate electrostatic, topological, algebraic, geometric, and features to train a learned model and then use the learned model to predict unknown properties for protein with available structures. This project focuses on electrostatic and topological features.
@@
The main goal of this project is to discover the hidden and useful information embedded in the protein structural data from the protein data bank in a simple and abstract way. The protein structure data and force field include bonded and non-bonded interactions with which the molecular simulation can be performed. We categorize these information into electrostatic, topological, algebraic, and geometric features. These four aspects are all important to the machine learning model. In this project, we put our focus on the electrostatic and topological features. The DNN based ML model is shown in Fig. 5, which uses the available protein structural data PDB, force field such as AMBER Perlman:1995, CHARMMBrooks:1983, AMOEBA Zhang:2018, etc. and known protein properties repositories such as PDBbind Database Su:2019, Protein pKa Database Pahari:2019, etc to mathematically generate features to train a learned model and then use the learned model to predict unknown properties for protein with available structures. In this project, we focus on the effectiveness of topological and electrostatic features using computed Coulomb energies and solvation energies as labels.
Mathematically, the surrogate models (Coulomb energies computed by treecode or solvation energies from the numerically solved PB model)
$E=f(\mathbf{X},\mathbf{Q})$ (15)
are approximated by the machine learning models
$\hat{E}=\hat{f}(F_{e},F_{t},\Theta)$ (16)
where $E$ and $\hat{E}$ are the energies, $\mathbf{X}$ and $\mathbf{Q}$ are the positions and charges of protein atoms, $\Theta$ are parameters in the ML model to be trained by data, $F_{e}=F_{e}(\mathbf{X},\mathbf{Q})$ and $F_{t}=F_{t}(\mathbf{X})$ are electrostatic features and topological features and their generators, which are non-trivial mappings from protein structure and electrostatics to features.
2.6 Specification of the ML Model
The model developed in this study is to to predict Coulomb and solvation energies using combined topological and electrostatic features. All target labels were standardized to zero mean and unit variance, and predictions were inverse-transformed to recover physical units (kcal/mol) for evaluation. As illustrated in Fig. 6, the proposed architecture consists of two branches. The first branch, shown in blue, is a one-dimensional convolutional neural network (CNN) designed to process the topological features. These features are passed through three consecutive convolution–pooling blocks (with 128, 64, and 64 filters; kernel size = 3; and activation functions tanh, ReLU, and tanh, respectively), followed by a fully connected layer with 64 units. The electrostatic features are processed through two fully connected hidden layers with 128 and 64 units, respectively (shown in orange). After independent processing in these two branches, their outputs are concatenated and passed through three additional fully connected layers with 32, 16, and 8 units (shown in purple). The final dense layer then produces the regression output for the predicted energy (shown in red). To improve generalization, dropout, $l_{2}$ regularization, and batch normalization were applied. The model was trained using the Adam optimizer (learning rate $1.0\times 10^{-4}$) with mean squared error as the loss function. The dataset was randomly divided into 80% training and 20% testing subsets. Model selection was performed using five-fold cross-validation on the training set, and the best-validated model was subsequently evaluated on the held-out test set. The model architecture and hyperparameters were selected based on empirical considerations and adopted from previously published studies Cang:2017. The complete hyperparameter configuration is provided in Table $S1$ of the Supporting Information.
2.7 Data
The datasets used in this study are sourced from the PDBbind database Liu:2014, a widely used benchmark for protein–ligand complexes. We categorize the data into two datasets as described below. For all structures, atomic van der Waals radii and partial charges are assigned using the AMBER ff14SB force field through the PDB2PQR software Dolinsky:2007.
[…]
Dataset 2. This dataset is constructed from the PDBbind v2020 release, comprising a general set of 14,127 complexes and a refined set of 5,316 complexes. Coulombic energies $E_{\mathrm{coul}}$ are computed for all proteins using a Cartesian treecode algorithm Li:2009. To ensure data quality, outliers are identified and removed using the IQR method, applying thresholds of [1,164.23, 338.04][-1,164.23,\;338.04] kcal/mol, where $Q_{1}=-600.88$ kcal/mol,  $Q_{3}=-225.31$ kcal/mol, $\text{IQR}=375.57$ kcal/mol. Following this filtering step, a total of 17,007 protein structures are retained for subsequent model training and evaluation.
(a)                (b)                (c)
Figure 7: Box-and-whisker plots showing the distributions of Coulombic energy $E_{\mathrm{coul}}$ for Dataset 1 in (a) and Dataset 2 in (b), and solvation energy $E_{\mathrm{solv}}$ for Dataset 1 in (c). Jittered points represent individual protein samples.
Figure 7 illustrates the distributions of Coulombic and solvation energies across the two datasets. The Coulombic energies in both datasets are concentrated within a relatively narrow negative range, whereas the solvation energies exhibit a substantially broader distribution with greater dispersion. This contrast reflects the fundamentally different physical scales and variability associated with Coulombic interactions and solvation effects.
2.8 Research Purposes and Metrics
We provide simulation results to evaluate the performance of the proposed model. These results are designed to address the following three research purposes:
Purpose 1: Training on a larger dataset leads to improved predictive performance.
Purpose 2: When electrostatic features are incorporated, model performance improves as the treecode level $L$, Taylor expansion order $p$, and the corresponding number of features $N_{f}$ increase.
Purpose 3: Combining electrostatic and topological features outperforms models that use only one type of feature.
Metrics: We use three common metrics to measure the performance of the model, as defined below. In all these metrics, we use $y_{i}$ to denote actual value, $\hat{y}_{i}$ for predicted value, $\bar{y}$ for mean of actual values, $\bar{\hat{y}}$ for mean of predicted values.
MAPE: The mean average percentage errors, which measures prediction accuracy as a percentage. This metric is not suitable if the actual values contain zero or values close to zero.
$\text{MAPE}=\frac{1}{n}\sum_{i=1}^{n}\left|\frac{y_{i}-\hat{y}_{i}}{y_{i}}\right|\times 100\%$ (17)
MSE: The mean-squared error, which measures the average squared difference between predicted and actual values.
$\text{MSE}=\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}$ (18)
$R^{2}$, the quantity indicating the proportion of variance in the dependent variable that is predictable from the independent variables.
$R^{2}=1-\frac{\text{SS}_{\text{res}}}{\text{SS}_{\text{tot}}}$ (19)